If the price of a slice of pizza rises from $2.50 to $3, and quantity demanded falls from 10,000 slices to 7,400 slices, using the formula for arc price elasticity

  What is the percentage change in quantity?


1
Expert's answer
2021-07-22T10:00:39-0400

Let's find the percentage change in quantity demanded:


"\%\Delta Q=\dfrac{Q_2-Q_1}{0.5\cdot(Q_2+Q_1)}\cdot100\%,""\%\Delta Q=\dfrac{7400-10000}{0.5\cdot(7400+10000)}\cdot100\%=-29.89\%."